OpenStudy (anonymous):

a+b+c=2[(a-b)+(a-c)] what does a=

OpenStudy (alfie):

a+b+c = 2[a-b+a-c] a+b+c = 2[2a-b-c] a+b+c= 4a-2b-2c a -4a = -2b-b-c -2c -3a = -3b -3c 3a = 3b + 3c a = (3b+3c)/3 a = b+c

OpenStudy (anonymous):

Thank you so much :)